# SuperMat: Physically Consistent PBR Material Estimation at Interactive Rates

### v3

@@ 3.3 UV Refinement Network
IntrinsicAnything [5] 20.7927 0.8721 0.1582 - - - - - - - - - 6.93
StableMaterial [23] 23.8992 0.9050 0.1032 18.7430 0.8321 0.2207 21.4639 0.9029 0.1168 22.8837 0.9072 0.0914 0.53
Ours w/o e2e 24.6155 0.8966 0.1102 19.8342 0.8576 0.2321 20.8168 0.8769 0.1391 24.3308 0.9125 0.0899 3.09
Ours w/o re-render 27.2149 0.9219 0.0862 22.9245 0.8474 0.1968 23.3083 0.9043 0.1122 26.7900 0.9341 0.0687 0.07
Ours 27.5034 0.9240 0.0845 23.4598 0.8598 0.1937 23.5158 0.9040 0.1114 27.1192 0.9366 0.0640 0.07
Table 1: Comparison of our method with others on the image space decomposition task. We highlight the best, second-best, and third-best results for each metric. ”Ours w/o e2e” denotes our method without the scheduler fix and perceptual loss for end-to-end training. ”Ours w/o re-render” denotes our method without the re-render loss. The first two methods are non-diffusion-based, while the middle four methods are diffusion-based.
We aim to apply SuperMat to decompose material for 3D objects with only RGB textures. Several previous works [29, 4, 45, 35] use techniques like SDS to bridge the gap between dimensions, but such methods are not applicable to our deterministic model, and suffer from slow computational speed and mode seeking issues. Therefore, we explore using backprojection to transfer 2D decomposition results from image space to UV space, enabling their application to 3D models. Notably, SuperMat is a single-view model, meaning that one inference can only decompose the visible region from a single viewpoint. To address this problem, we propose estimating materials from multiple viewpoints and integrating them. With multi-view decomposition results from SuperMat, the UV refinement network aims at inpainting missing areas, eliminating discrepancies between different views, and improving the overall quality of the generated materials.
Backprojection and UV Blending. For each viewpoint, we backproject SuperMat’s decomposition results from image space onto the model’s UV texture map, obtaining the partial albedo UV map $C_{d}^{i}$, RM UV map $C_{rm}^{i}$, and mask UV map $M^{i}$ for viewpoint $i$. We then blend the backprojection results from all viewpoints in an averaged manner. For albedo, this can be formulated as:
@@
The RM follows the same computation method. The blended partial maps $C_{d}$, $C_{rm}$, and the partial mask $M$ are then input into the subsequent UV refinement one-step model.
UV Refinement One-Step Model. Due to the limited number of viewpoints, the blended partial maps often have missing areas. Additionally, slight variances may exist between SuperMat’s results from different views. To address these issues, we leverage diffusion models for UV refinement: we fine-tune Stable Diffusion into a UV refinement model, which takes a flawed partial material UV map, a mask UV map, and a position UV map as inputs, and outputs completed, variance-corrected, and refined material UV maps. Specifically, the material maps are encoded by a VAE and concatenated with downsampled mask and position maps, and then processed by UNet with an expanded 8-channel conv_in. We also modify the model’s scheduler to support single-step inference and enable end-to-end training, resulting a deterministic one-step model.
Material Decomposition Pipeline for 3D Objects. As shown in Fig. 3, with the fine-tuned UV refinement network, we propose a complete pipeline for the material decomposition of 3D objects. Given an object with only RGB texture, we first render it from six orthogonal viewpoints. The rendered results are then batch-processed by SuperMat for material decomposition. These materials, including albedo and RM, are backprojected into UV space and blended. The blended partial maps are then refined by the UV refinement one-step model and the final material UV maps are generated. This process requires only three inference steps—one for SuperMat, one for refining albedo, and one for refining RM—allowing our pipeline to complete the material decomposition of a 3D object in just 3 seconds.
@@ 4 Experiments
Figure 4: Comparison of our method with others on albedo estimation results.
@@
Figure 8: Results of decomposition for 3D objects. We show the comparison between the original and decomposed models under the same environment map, along with the relighting outcomes of new models.
Figure 9: Qualitative ablation results on UV refinement network.
Quantitative Comparison of Image Space Decomposition. We randomly select 64 unseen items each from the Objaverse [8] and BlenderVault [23] datasets. Our method is tested against 6 baseline approaches on these 128 objects. For each object, the model processes 18 inputs corresponding to 3 lighting conditions and 6 viewpoints, predicting 3 (or 1) types of materials and generating 54 (or 18) texture maps in total. All outputs are used to compute metrics, with the final averaged results shown in Tab.1. We use PSNR, SSIM, and LPIPS metrics to evaluate albedo, metallic, and roughness. Additionally, for methods that provide all material types, we compute metrics on the rendering under relighting conditions using the ground truth normals and positions and the predicted values. The results demonstrate that our approach significantly outperforms the state-of-the-art in image space decomposition.
Beyond image quality, we compare inference speeds by calculating the total time required for each method to process the test set, normalized by the dataset size and number of predicted material types. All tests are conducted with a batch size of 1. For consistency in speed comparison, StableMaterial and IID, which typically average over 10 predictions per input, are modified to perform a single prediction. The “Time(s)” column in Tab.1 reflects the time required by each method to generate a single material prediction on an NVIDIA A100 GPU with a batch size of 1. Traditional, non-diffusion-based methods like Derender3D and IIR generally show faster speeds. While StableMaterial attempts to predict multiple materials simultaneously by adding extra channels to conv_in and conv_out, its lack of structural design adjustments limits prediction quality and requires multi-step denoising, resulting in slower performance than traditional methods. In contrast, our approach achieves high-quality material decomposition with speeds comparable to traditional methods.
@@
Qualitative Comparison of Image Space Decomposition. Fig.4 compares albedo prediction results across various cases, including artist-designed objects and real-world scans, showcasing the performance of our method against other approaches. SuperMat demonstrates superior ability in removing highlights and shadows, accurately capturing the intrinsic color, even on challenging metallic surfaces. In Fig.5, we present metallic and roughness predictions from our method alongside other baselines. SuperMat outperforms the others with well-defined boundaries between metallic and non-metallic regions and achieves consistency within material types while preserving high-frequency details. Moreover, our results closely align with physical realism, addressing a significant challenge that previous methods struggled to overcome. Results with a wider variety of inputs will be presented in the supplementary materials.
@@
Results of UV Refinement and Material Estimation for 3D Objects. Fig.6 illustrates the effectiveness of the UV refinement one-step model in inpainting missing areas and harmonizing viewpoint variances in SuperMat’s outputs. Fig.8 displays the results from our material estimation pipeline for 3D models. The first column shows the rendering of the input models with only RGB texture, where all surfaces exhibit uniform reflectance behavior. After applying our method, the decomposed materials enable physically-based rendering with surface-specific reflectance properties, as demonstrated in the second and third columns under different lighting conditions.
@@
Ablation Studies. In Tab.1, we present ablation studies on two core components of SuperMat: end-to-end (e2e) training and re-render loss. E2E with perceptual loss significantly enhances model performance and efficiency. In fact, within the e2e framework, we experiment with three types of loss functions, with perceptual loss achieving the best results. This suggests that direct supervision of the fully denoised predicted material maps, rather than latent noise, plays a critical role in enhancing the model’s capability for this specific task. Detailed findings are available in the supplementary materials. The addition of re-render loss, facilitated by the e2e framework, further improves prediction quality, particularly under relighting conditions. Qualitative results of these ablations are shown in Fig. 7. The ‘w/o e2e’ setting, where the model is trained without perceptual and re-render losses, limits its capacity to remove lighting effects effectively. In contrast, training ‘w/o re-render’ loss leads to minor inaccuracies in fine material details. Note that in the ’w/o e2e’ setting, there are slight differences in the implementation of the UNet. Detailed explanations will be provided in the supplementary materials.
@@ 6 Additional Details
Implementation Details. We report some implementation details as follows: 1) In the UV refinement one-step model, the input channels are expanded to 8, with the weights of the additional 4 channels initialized to 0. 2) In the ablation experiment, the structure of SuperMat under the “w/o e2e” setting is slightly different. Without single-step inference, we can only train the diffusion model in a denoising task manner, meaning that during the single-step denoising process, the latents that the UNet receives are not encoded from a clean rendered image, but rather a noisy albedo and noisy RM. Therefore, on top of SuperMat, we additionally replicate the conv_in and the first DownBlock as independent parts of structural expert branches to map the inputs from two different domains to similar distributions. These are then fused in an averaged manner before being fed into the shared modules. 3) In the re-render loss implementation, each time we perform relighting, we randomly select a lighting condition from a set of 716 environment maps, covering nearly all possible lighting scenarios.
Training Details. Both models are fine-tuned from Stable Diffusion 2.1. We train SuperMat using the AdamW optimizer with a learning rate of $51e-$ 5 on 8 NVIDIA A800 (80GB) GPUs, with a batch size of 16, for a total of 115 epochs. The UV refinement one-step model is trained with the AdamW optimizer at a learning rate of $52e-$ 5, also on 8 NVIDIA A800 (80GB) GPUs, with a batch size of 16, for 40 epochs. The images are resized to resolutions of $512\times 512$ and $1024\times 1024$, used for SuperMat and the UV refinement one-step model, respectively.
Training Datasets. In this paper, we collect 3D objects from Objaverse [8]. For each object, we render material maps, normal maps, and position maps from six orthogonal viewpoints and generate PBR results under three different environment lighting conditions from the same viewpoints. We then filter high-quality data using methods such as DINO score considerations [28], resulting in a dataset of 100k objects for the image space decomposition task. Following the approach outlined in Sec. 3.3, we further refine this data for the UV refinement network by selecting cases where the missing area in blended partial UV maps is under $50\%$. This process results in a final dataset of 40k objects for training the refinement one-step model.
@@ 7 Additional Visualizations
In Figures 10111213, we present additional results of SuperMat on the Objaverse and BlenderVault test datasets, covering both artist-designed and real-world scanned objects. Fig. 14 highlights more decomposition results of real-world objects extracted from captured images. Finally, Fig. 15 illustrates more decomposition results on 3D objects.

### v4

@@ 3.2 UV Refinement Network
IntrinsicAnything [6] 20.94 0.8788 0.1520 - - - - - - - - - 6.93
StableMaterial [30] 23.44 0.9019 0.1076 20.29 0.8164 0.2459 21.01 0.8880 0.1327 22.56 0.9043 0.0979 0.53
StableMaterialMV [30] 24.56 0.9091 0.0946 17.28 0.7886 0.2829 20.86 0.8838 0.1286 23.14 0.9090 0.0933 0.55
SuperMat w/o e2e 24.26 0.8964 0.1095 20.79 0.8444 0.2826 20.81 0.8621 0.1587 23.90 0.9085 0.0964 3.09
SuperMat w/o re-render 26.70 0.9236 0.0906 24.54 0.8259 0.2263 23.52 0.9016 0.1362 26.41 0.9378 0.0702 0.07
SuperMat 27.68 0.9290 0.0820 25.48 0.8685 0.2221 24.25 0.9057 0.1281 27.66 0.9432 0.0612 0.07
SuperMatMV 27.56 0.9289 0.0839 26.11 0.8749 0.2195 24.84 0.9076 0.1208 27.64 0.9431 0.0628 0.09
Table 1: Comparison of our method with others on the image space decomposition task. We highlight the best, second-best, and third-best results for each metric. ”SuperMat w/o e2e” denotes our method without the scheduler fix and perceptual loss for end-to-end training. ”SuperMat w/o re-render” denotes our method without the re-render loss. The first two methods are non-diffusion-based, while the middle five methods are diffusion-based.
We aim to apply SuperMat or SuperMatMV to decompose material for 3D objects with only RGB textures. Several previous works [38, 5, 65, 48] use techniques like SDS to bridge the gap between dimensions, but such methods are not applicable to our deterministic model, and suffer from slow computational speed and mode seeking issues. Therefore, we explore using backprojection to transfer 2D decomposition results from image space to UV space, enabling their application to 3D models. Notably, both SuperMat and SuperMatMV are finite-view models, meaning that one inference can only decompose the visible region from a specific set of viewpoints. To address this problem, we propose estimating materials from multiple viewpoints and integrating them. With multi-view decomposition results from SuperMatMV, the UV refinement network aims to paint missing areas and improve the quality of the generated materials.
Backprojection and UV Blending. For each viewpoint, we backproject SuperMatMV’s decomposition results from image space onto the model’s UV texture map, obtaining the partial albedo UV map $C_{d}^{i}$, RM UV map $C_{rm}^{i}$, and mask UV map $M^{i}$ for viewpoint $i$. We then blend the backprojection results from all viewpoints in an averaged manner. For albedo, this can be formulated as:
@@
The RM follows the same computation method. The blended partial maps $C_{d}$, $C_{rm}$, and the partial mask $M$ are then input into the UV refinement one-step model.
UV Refinement One-Step Model. Due to the limited number of viewpoints, the blended partial maps often have missing areas. To address this issue, we leverage diffusion models for UV refinement: we fine-tune Stable Diffusion into a UV refinement model, which takes a flawed partial material UV map, a mask UV map, and a position UV map as inputs, and outputs completed and refined material UV maps. Specifically, the material maps are encoded by a VAE and concatenated with downsampled mask and position maps, and then processed by UNet with an expanded 8-channel conv_in. We also modify the model’s scheduler to support single-step inference and enable end-to-end training, resulting in a deterministic one-step model.
Material Decomposition Pipeline for 3D Objects. As shown in Fig. 3, with the fine-tuned UV refinement network, we propose a complete pipeline for the material decomposition of 3D objects. Given an object with only RGB texture, we first render it from six orthogonal viewpoints. The rendered results are then batch-processed by SuperMatMV for viewpoint-consistent material decomposition. These materials, including albedo and RM, are backprojected into UV space and blended. The blended partial maps are then refined by the UV refinement one-step model and the final material UV maps are generated. This process requires only three inference steps—one for SuperMatMV, one for refining albedo, and one for refining RM—allowing our pipeline to complete the material decomposition of a 3D object in just 3 seconds.
@@ 4 Experiments
Figure 4: Comparison of our method with others on albedo estimation results. Each pair of rows presents results of the same object under identical lighting conditions from different viewpoints. This organizational format is also followed in other qualitative comparison figures.
@@
Training Datasets. In this paper, we collect 3D objects from Objaverse [9] and Digital Twin Catalog (DTC) [14] datasets. For each object, we render material maps, normal maps, and position maps from 6 orthogonal viewpoints and generate PBR results under 16 different environment lighting conditions from the same viewpoints. We then filter high-quality data using methods such as DINO score considerations [37], resulting in a dataset of 75k objects that combines both synthetic and real-captured data for the image space decomposition task. Following the approach outlined in Sec. 3.2, we further refine this data for the UV refinement network by selecting cases where the missing area in blended partial UV maps is under $50\%$. This process results in a final dataset of 40k objects for training the refinement one-step model.
Quantitative Comparison of Image Space Decomposition. We randomly select 64 unseen items each from the Objaverse, DTC and BlenderVault [30] datasets. Objaverse and BlenderVault primarily contain synthetic data, whereas DTC consists of scanned objects, offering a source of real-world data. Our method is tested against 7 baseline approaches on these 192 objects. For each object, the model processes 18 inputs corresponding to 3 lighting conditions and 6 viewpoints, predicting 3 (or 1) types of materials and generating 54 (or 18) texture maps in total. All outputs are used to compute metrics, with the final averaged results shown in Tab.1. We use PSNR, SSIM, and LPIPS metrics to evaluate albedo, metallic, and roughness. Additionally, for methods that provide all material types, we compute metrics on the rendering under relighting conditions using the ground truth normals and positions and the predicted values. The results demonstrate that our approach significantly outperforms the SOTA in image space decomposition.
Beyond image quality, we compare inference speeds by calculating the total time required for each method to process the test set, normalized by the dataset size and number of predicted material types. All tests are conducted with a batch size of 1. For consistency in speed comparison, IID, StableMaterial and StableMaterialMV, which typically average over 10 predictions per input, are modified to perform a single prediction. The “Time(s)” column in Tab.1 reflects the time required by each method to generate a single material prediction on an NVIDIA A100 GPU with a batch size of 1. Traditional, non-diffusion-based methods like Derender3D and IIR generally show faster speeds. While StableMaterial attempts to predict multiple materials simultaneously by adding extra channels to conv_in and conv_out, its lack of structural design adjustments limits prediction quality and requires multi-step denoising, resulting in slower performance than traditional methods. In contrast, our approach achieves high-quality material decomposition with speeds comparable to traditional methods.
@@
Figure 8: Results of decomposition for 3D objects. We compare the original and decomposed models under the same illumination and show the relighting outcomes of the new models.
@@
Qualitative Comparison of Image Space Decomposition. Fig.4 compares albedo prediction results across various cases, including artist-designed objects and real-world scans, showcasing the performance of our method against other approaches. SuperMat demonstrates superior ability in removing highlights and shadows, accurately capturing the intrinsic color, even on challenging metallic surfaces. In Fig.5, we present metallic and roughness predictions from our method alongside other baselines. SuperMat outperforms the others with well-defined boundaries between metallic and non-metallic regions and achieves consistency within material types while preserving high-frequency details. Moreover, our results closely align with physical realism, addressing a significant challenge that previous methods struggled to overcome. Additionally, SuperMatMV demonstrates stronger consistency across different views in all material predictions, and is still able to generate reasonable decomposition results for inputs from challenging viewpoints with sparse relevant information. More results will be presented in the supplementary materials.
@@
Results of UV Refinement and Material Estimation for 3D Objects. Fig.6 illustrates the effectiveness of the UV refinement one-step model in inpainting missing areas in SuperMatMV’s outputs. Fig.8 displays the results from our material estimation pipeline for 3D models. The first column shows the rendering of the input models with only RGB texture, where all surfaces exhibit uniform reflectance behavior. After applying our method, the decomposed materials enable physically-based rendering with surface-specific reflectance properties, as demonstrated in the second and third columns under different lighting conditions.
@@
Ablation Studies. In Tab.1, we present ablation studies on two core components of SuperMat: end-to-end (e2e) training and re-render loss. E2E with perceptual loss significantly enhances model performance and efficiency. In fact, within the e2e framework, we experiment with three types of loss functions, with perceptual loss achieving the best results. This suggests that direct supervision of the fully denoised predicted material maps, rather than latent noise, plays a critical role in enhancing the model’s capability for this specific task. Detailed findings are available in the supplementary materials. The addition of re-render loss, facilitated by the e2e framework, further improves prediction quality, particularly under relighting conditions. Qualitative results of these ablations are shown in Fig. 7. The ‘w/o e2e’ setting, where the model is trained without perceptual and re-render losses, limits its capacity to remove lighting effects effectively. In contrast, training ‘w/o re-render’ loss leads to minor inaccuracies in fine material details.
@@ 5 Conclusions
Figure 9: Qualitative ablation results on UV refinement network.
@@
Albedo Metallic Roughness Relighting
L1 26.8102 0.9164 0.0956 22.4801 0.8889 0.1803 23.5855 0.9109 0.1194 26.8610 0.9374 0.0672
SSI 24.2359 0.9063 0.1033 22.2521 0.8701 0.1814 23.7918 0.9139 0.1165 24.7511 0.9266 0.0732
Perceptual 27.0082 0.9151 0.0949 22.8702 0.8669 0.1760 24.1452 0.9145 0.1156 27.2484 0.9374 0.0650
Training Data Type Albedo Metallic Roughness
MSE↓ PSNR↑ SSIM↑ LPIPS↓ MSE↓ PSNR↑ SSIM↑ LPIPS↓ MSE↓ PSNR↑ SSIM↑ LPIPS↓
Synthetic 0.0027 27.41 0.9195 0.0885 0.0120 23.80 0.8767 0.1692 0.0084 24.32 0.9152 0.1118
Real-captured 0.0061 24.50 0.8982 0.1049 0.0350 19.40 0.8464 0.2043 0.0113 22.32 0.9075 0.1165
Synthetic+Real-captured 0.0025 27.58 0.9205 0.0867 0.0111 23.77 0.8787 0.1696 0.0074 24.63 0.9154 0.1115
Table 3: Ablation results on training data types in the image space decomposition task. We highlight the best results for each metric.
Albedo RM
PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓
Input 13.56 0.5499 0.4227 12.52 0.5484 0.5245
Refined 23.92 0.7537 0.2792 27.31 0.8460 0.1178
Table 4: Quantitative evaluation on UV maps before and after refinement.
BlenderVault StanfordORB
PSNR↑ SSIM↑ LPIPS↓ PSNR↑ SSIM↑ LPIPS↓
Derender3D 22.52 0.9151 0.1023 17.03 0.8241 0.2012
IIR 21.10 0.9005 0.1046 15.83 0.7944 0.2005
IID 22.10 0.9095 0.0989 16.21 0.8000 0.2059
RGB→X 22.23 0.9115 0.0987 17.23 0.8272 0.1849
IA 23.33 0.9115 0.1047 17.98 0.8286 0.1875
SM 23.36 0.9176 0.0909 16.94 0.8174 0.1801
SMMV 23.57 0.9175 0.0887 17.10 0.8189 0.1785
Ours 27.00 0.9463 0.0629 24.99 0.9281 0.0962
Table 5: Quantitative comparison of novel view synthesis relighting. We highlight the best, second-best, and third-best results for each metric. Here “IA”, “SM”, “SMMV” represents “IntrinsicAnything”, “StableMaterial”, “StableMaterialMV” respectively.
6 Physically Based Rendering (PBR) Model
We use the Cook-Torrance Bidirectional Reflectance Distribution Function (BRDF) [8] based on microfacet theory to define the materials and establish the rendering model. Following [5], for a point with coordinates $p\in\mathbb{R}^{3}$, albedo $a\in\mathbb{R}^{3}$, metallic $m\in\mathbb{R}$, roughness $r\in\mathbb{R}$, and surface orientation $n\in\mathbb{R}^{3}$, the PBR result $L$ observed from viewpoint $c\in\mathbb{R}^{3}$ is given by:
@@ 7 Additional Details
Implementation Details. We report some implementation details as follows: 1) In the UV refinement one-step model, the input channels are expanded to 8, with the weights of the additional 4 channels initialized to 0. 2) In the ablation experiment, the structure of SuperMat under the “w/o e2e” setting is slightly different. Without single-step inference, we can only train the diffusion model in a denoising task manner, meaning that during the single-step denoising process, the latents that the UNet receives are not encoded from a clean rendered image, but rather a noisy albedo and noisy RM. Therefore, on top of SuperMat, we additionally replicate the conv_in and the first DownBlock as independent parts of structural expert branches to map the inputs from two different domains to similar distributions. These are then fused in an averaged manner before being fed into the shared modules. 3) In the re-render loss implementation, each time we perform relighting, we randomly select a lighting condition from a set of 50 environment maps, covering nearly all possible lighting scenarios.
Training Details. Both the SuperMat and the UV refinement model are fine-tuned from Stable Diffusion 2.1, while SuperMatMV is built upon SuperMat. We train SuperMat using the AdamW optimizer with a learning rate of 2e52e-5 on 8 NVIDIA A800 (80GB) GPUs, with a batch size of 32, for a total of 30 epochs. The UV refinement one-step model is trained with the AdamW optimizer at a learning rate of 2e52e-5, also on 8 NVIDIA A800 (80GB) GPUs, with a batch size of 16, for 40 epochs. The images are resized to resolutions of $512\times 512$ and $1024\times 1024$, used for SuperMat and the UV refinement one-step model, respectively. The training setup for SuperMatMV mirrors that of SuperMat, except that SuperMatMV is trained for only 3 epochs using a batch size of 8 for 6-view images.
@@ 8 Additional Visualizations
In Figures 111213141516, we present additional results of SuperMat on the Objaverse [9], BlenderVault [30] and DTC [14] test datasets, covering both artist-designed and real-world scanned objects. Fig. 17 illustrates more decomposition results on 3D objects.
9 Additional Experiments
Novel View Synthesis Relighting. We validate our material decomposition pipeline for 3D objects by relighting the decomposed objects under novel lighting and viewpoints and comparing the results with ground-truth relit object images. We select all 14 objects from the StanfordORB dataset [27] and 14 randomly chosen unseen objects from the BlendVault dataset [30], applying appropriate lighting conditions before performing material decomposition. We adapt baseline methods to incorporate the same UV backprojection and blending framework as our approach, allowing them to handle 3D object decomposition. We then render the decomposed models from 60 novel viewpoints under three different environment maps. The quantitative comparison results are presented in Tab. 5, while qualitative results are shown in Fig. 10. Our method significantly outperforms others in both texture completeness and the physical consistency of materials, demonstrating its effectiveness in decomposing high-quality materials for 3D objects.
Comparison between Different Loss Functions. Without re-render loss, we experiment with three types of loss functions in the end-to-end framework: L1 loss, shift and scale invariant (SSI) loss [40], and perceptual loss. The performance of models trained with these loss designs is evaluated on the same test dataset in Objaverse [9], and the results are shown in Tab. 2. Among these losses, the perceptual loss, which is ultimately adopted by SuperMat, demonstrates the best performance.
Quantitative Validation of UV Refinement. Tab. 4 presents the quantitative evaluation results of the UV maps generated and blended by SuperMatMV before and after refinement, validating the effectiveness of the UV refinement one-step model.
Ablation on the Training Dataset. For the image space decomposition task, we conduct an ablation study on the composition of the training dataset, testing the SuperMat’s performance when trained solely on synthetic data or only on real-captured data. We randomly select 64 objects containing both data types from the training set, totaling 6144 inputs, which are unseen during training but used to evaluate the model’s estimations of 3 material types with MSE, PSNR, SSIM, and LPIPS metrics. The results of the experiment are shown in Tab.3, indicating that data diversity helps improve the model’s generalization ability.
Detailed Quantitative Comparison. We show the detailed quantitative comparison of the image space decomposition task in Tables 678 and  9.
10 Limitations
Although our method enables fast and physically consistent material decomposition for both images and 3D objects, it still has several limitations. First, converting the diffusion model into a deterministic model significantly improves performance but also sacrifices the benefits of the diffusion process. For instance, in cases with high-frequency details, SuperMat’s results may lack sharpness. Additionally, as a deterministic model, SuperMat has reduced flexibility—once an estimation error occurs, rerunning the process will always yield the same incorrect result. Furthermore, due to the constraints of the chosen BRDF model, SuperMat faces challenges in handling transparent materials, multi-layered surfaces, and highly reflective objects.
Figure 10: Qualitative comparison of novel view synthesis relighting.